OpenStudy (anonymous):

PLEASE HELP Explain the meaning of absolute value of a complex number. What is the absolute value of –2 + 3i? Can someone help me through step by step?

OpenStudy (anonymous):

Wasn't this just answered about 15 minutes ago?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I just answered the same question about fifteen minutes ago from someone else. Can you check the log of questions from 15 minutes ago?

OpenStudy (anonymous):

well do you think you can help with the first part?

OpenStudy (anonymous):

absolute value of -2+3i is sqrt((-2)^2 + (3)^2) = sqrt (4+9) = sqrt(13). It represents the distance from the point -2+3i to the origin. So both parts are answered.